# Optimization of Quantum-Repeater Networks using Stochastic Automatic Differentiation

### v1

@@ I Introduction
where the variance of $X^{\prime}(p)$ is bounded. Roughly speaking, stochastic AD achieves this feat by not estimating the value of $\mathbb{E}[X(p+\epsilon)]$ and $\mathbb{E}[X(p)]$ independently, as one would do in finite differences, but instead uses shared randomness between the two evaluations. Moreover, the programs created by stochastic AD are composable, meaning that when a random variable $X(p)$ is a function of random variables $X_{1}(p)$ and $X_{2}(p)$, a program that samples $X^{\prime}(p)$ can be constructed from programs that sample $X^{\prime}_{1}(p)$ and $X^{\prime}_{2}(p)$. This enables the technique to be used as a form of automatic differentiation (AD). While AD is a well established technique, traditional versions of AD do not work in the context of discrete randomness, which is crucial for simulations of quantum networks. We will further explain AD and stochastic AD in Section II.
In this paper, we apply stochastic AD to obtain derivatives from Monte Carlo simulations of quantum-repeater networks and demonstrate how this enables their optimization and analysis. First, in Section II, we further explain what (stochastic) AD is and how it works. We remark though that understanding stochastic AD is not needed to understand the results obtained using stochastic AD. In Section III, we explain the quantum-network model that is employed throughout the rest of the paper. Then, in Section IV, we give four different demonstrations of the usefulness of stochastic AD for quantum networks. In Section IV.1, we explicitly compare stochastic AD to finite differences. In Section IV.2, we use gradient descent enabled by stochastic AD to optimize rate-fidelity tradeoffs on the elementary links of a quantum-repeater chain. In Section IV.3, we perform a sensitivity analysis with respect to the coherence times at the different nodes in the same repeater chain. Finally, we present our most extensive result in Section IV.4, where we investigate the optimal placement of quantum repeaters in a 2D plane.
@@ III Model for quantum repeaters
Each pair of neighboring nodes in the chain is connected by an elementary link that can be used to generate elementary-link entanglement. In a chain of $N$ nodes (of which $N-2$ are repeaters), there are $N-1$ elementary links. Generating elementary-link entanglement is a sequence of probabilistic attempts. Each attempt at elementary link $i$ succeeds only with probability $p_{i}$ and takes an amount of time $\Delta t_{i}$. We assume that the attempt duration is limited by the transmission of optical pulses through fiber, giving
@@
where $L_{i}$ is the length of elementary link $i$ and $c$ is the speed of light in fiber, which we assume to equal $200,000$ km/s (approximately two thirds the speed of light in vacuum). When an attempt is successful, at the end of the corresponding time slot, the two-qubit state $\rho_{i}$ is created between the two nodes. Ideally this is the Bell state, $\rho_{i}=\ket{\phi^{+}}\bra{\phi^{+}}$ | with |$\ket{\phi^{+}}=\tfrac{1}{\sqrt{2}}\left(\ket{00}+\ket{11}\right)$, but more realistically there will be noise and $\rho_{i}$ will be some mixed state instead. We here assume all elementary-link states are Werner states [81, 17] of the form
$\rho_{i}=w_{i}\ket{\phi^{+}}\bra{\phi^{+}}+(1-w_{i})\frac{\mathbb{1}}{4},$ (7)
@@
$F_{i}\equiv\braket{\phi^{+}}{\rho_{i}}{\phi^{+}}=\frac{1+3w_{i}}{4}.$ (8)
@@ IV.2 Optimizing a repeater chain
$\sqrt{1-\alpha}\ket{\text{dark}}+\sqrt{\alpha}\ket{\text{bright}},$ (16)
where $\alpha$ is the bright-state parameter. After that the emitter is excited, resulting in photon emission only if the emitter is in the bright state, thereby creating entanglement between the emitter and the absence/presence of a photon. Photons emitted by both nodes are then directed to a midpoint station where an entangling measurement is performed that effectively counts the total number of photons that arrive. If only a single photon is measured, in the absence of photon loss, this projects the emitters on the state $\tfrac{1}{\sqrt{2}}(\ket{\text{dark, dark}}+\ket{\text{bright, bright}}$. However, when two photons are emitted but one is lost, this will also result in only a single photon being measured, and hence the entanglement will be noisy. The noise can be suppressed by making $\alpha$ small, but this comes at the cost of a reduced success rate. Hence, it tunes a tradeoff between rate and fidelity.
Using the repeater-chain model introduced in Section III, we here address the problem of choosing the bright-state parameters for each of the $N-1$ elementary links such that the secret-key rate, defined in Eq. (11), is optimized. A simplified version of this optimization was phrased as a mathematical program by Vardoyan et al. [25] that allows for convex optimization for a specific class of utility functions, a class that was extended further in [94]. Optimization of bright-state parameters under more complicated conditions was performed using genetic algorithms in [34, 38]. Here we follow Vardoyan et al. by making the typical simplifying assumption that each elementary link produces Werner states. In particular, when $\alpha_{i}$ is the bright-state parameter of elementary link $i$, states on that link are created with Werner parameter
@@ IV.4 Repeater placement
In Figure 4, we show the optimal value of $\text{SKR}_{\text{min}}$ that was found for $N=0,1,2,...,40$ for fixed $D=300$ km.
Figure 4: Dependence of the optimized network utility ($\text{SKR}_{\text{min}}$, defined in Eq. (19)) on the number of repeaters ($N$). End nodes form a square with edges of length $D=300$ km. Shown are the optimization results, a fit function as described by Eq. (22), the minimal number of repeaters ($N_{\text{min}}$, defined in Eq. (21)), the sufficient number of repeaters ($N_{\text{suff, }\delta}$, defined in Eq. (23), with $\delta=0.05$), and the best secret-key rate ($\text{SKR}_{\text{best}}$, defined in Eq. (24)). The error bars in the data are smaller than the marker size and have been suppressed.
@@ V Conclusion
Two open-source Julia packages were developed to make the investigations reported on in this paper possible, QuantumNetworkRecipes.jl [86] and RepeaterPlacement.jl [119]. All the code used for this work has been bundled in a GitHub repository [123], which has been archived at DOI:10.5281/zenodo.14595893.
@@
The data presented in this work has been made available at DOI:10.5281/zenodo.14594606.
@@ Appendix D Simulated annealing
In order to optimize repeater placement as discussed in Section IV.4, we have used simulated annealing. There is a temperature parameter that starts at a high value and is decreased every epoch. Each of the optimizations reported here consists of 25 epochs, with 15 iterations per epoch. For each iteration in epoch $i=1,2,...,25$, the used temperature is $t=t_{0}*0.8^{i}$. The value of $t_{0}$ at $D=300$ km is either 5, 10, or 15, and for other values of $D$, these values are changed in a way that is directly proportional to $D$. As reported in Section IV.4, the temperature is used as the learning rate of the gradient-descent algorithm; this is the reason why we set $t_{0}$ proportional to $D$, as the step size with which the repeaters are displaced should scale with the distances between the end nodes, as otherwise the algorithm will place repeater nodes far outside of the convex hull of the end nodes for small values of $D$ while they will not be varied significantly enough for large values of $D$. Apart from being used as the learning rate, the temperature affects the optimization in the following ways:
@@ Appendix E More repeater-placement results
In this appendix, we show some additional solutions for the repeater-placement optimization discussed in Section IV.4. In Figure 3, we have shown the best repeater placements we have found for $N=4,5,6,7$ and $D=300$ km. Here, in Figure 9, we give some examples of how solutions look for larger values of $N$.
@@
We observe that as the number of repeaters grows, the solutions become more irregular and less symmetric. This is likely the case because the optimization problem is more difficult than for a smaller number of repeaters and hence the optimization has not had the chance to converge as well within the 25 epochs of 15 iterations each. Better results would require more computational resources or more efficient optimizations (as discussed in Appendix C, pathfinding is currently the main bottleneck in the optimizations, especially when the number of repeaters is large). Moreover, we observe that as $N$ increases, an increasing number of repeaters is not part of any of the optimal paths between the end nodes. Therefore, the network utility ($\text{SKR}_{\text{min}}$, as defined in Eq. (19)) would not decrease in case those repeaters were removed. This is in line with our observation that the utility stops growing with the number of repeaters for large values of $N$: at some point, adding more repeaters does not allow one to increase the network utility, as introducing them to the paths would reduce the end-to-end fidelity for only a small gain in entangling rate.
In Figure 4 we have shown the dependence of the optimized network utility function $\text{SKR}_{\text{min}}$ on the number of repeaters $N$ for $D=300$ km. Here, in Figure 10, we show what this dependence looks like for a few different values of $D$.
Figure 10: Dependence of the guaranteed minimal quality of service, $\text{SKR}_{\text{min}}$ as defined in Eq. (19), on $N$ for $D=50$ km and $D=500$ km. Shown are the optimization results, a fit function as described by Eq. (22), the minimal number of repeaters ($N_{\text{min}}$, defined in Eq. (21)), the sufficient number of repeaters ($N_{\text{suff, }\delta}$, defined in Eq. (23), with $\delta=0.05$), and the best secret-key rate ($\text{SKR}_{\text{best}}$, defined in Eq. (24)). The error bars in the data are smaller than the marker size and have been suppressed.
For $D=500$ km, while it is clearly visible that the rate at which $\text{SKR}_{\text{min}}$ increases with $N$ slows down for larger values of $N$, it still appears to be growing at the largest value, $N=40$. For this reason, the estimates of $N_{\text{suff},\delta}$ and $\text{SKR}_{\text{best}}$ are perhaps a bit less reliable than those for smaller values of $D$, as a significant part of the function may still be hidden beyond $N=40$. Obtaining significant data for $D>500$ km would probably require solving the optimization for larger values of $N$, which may become computationally expensive.

### v2

@@ I Introduction
where the variance of $X^{\prime}(p)$ is bounded. Roughly speaking, stochastic AD achieves this feat by not estimating the value of $\mathbb{E}[X(p+\epsilon)]$ and $\mathbb{E}[X(p)]$ independently, as one would do in finite differences, but instead uses shared randomness between the two evaluations. Moreover, the programs created by stochastic AD are composable, meaning that when a random variable $X(p)$ is a function of random variables $X_{1}(p)$ and $X_{2}(p)$, a program that samples $X^{\prime}(p)$ can be constructed from programs that sample $X^{\prime}_{1}(p)$ and $X^{\prime}_{2}(p)$. This enables the technique to be used as a form of automatic differentiation (AD). While AD is a well established technique, traditional versions of AD either do not work or are inefficient in the context of discrete randomness, which is crucial for simulations of quantum networks. We will further explain AD and stochastic AD in Section II.
Being able to calculate derivatives for discrete random variables is particularly important in quantum mechanics, as any measurement in a finite-dimensional Hilbert space is discretely random. Using AD to differentiate through discretely random quantum evolution has been done before in the context of model-aware reinforcement learning, using the REINFORCE gradient estimator [74]. Examples include reinforcement learning for optimal control of quantum systems [77] and for quantum metrology [78, 79]. While REINFORCE is general purpose and unbiased, it is known for having a large variance, which often necessitates using tricks for variance reduction. In contrast, stochastic AD, which we use in this paper, does not require such tricks and provides a much smaller variance. A comparison between REINFORCE and stochastic AD for a discrete random walk is included in [76], demonstrating that stochastic AD provides a significantly smaller variance even when variance-reduction tricks are used for REINFORCE. To the best of our knowledge, AD has never been used in any form to differentiate through the dynamics of quantum-repeater networks. In this paper, we apply stochastic AD to obtain derivatives from Monte Carlo simulations of quantum-repeater networks and demonstrate how this enables their optimization and analysis.
This paper is structured as follows. First, in Section II, we further explain what (stochastic) AD is and how it works. We remark though that understanding stochastic AD is not needed to understand the results obtained using stochastic AD. In Section III, we explain the quantum-network model that is employed throughout the rest of the paper. Then, in Section IV, we give four different demonstrations of the usefulness of stochastic AD for quantum networks. In Section IV.1, we explicitly compare stochastic AD to finite differences. In Section IV.2, we use gradient descent enabled by stochastic AD to optimize rate-fidelity tradeoffs on the elementary links of a quantum-repeater chain. In Section IV.3, we perform a sensitivity analysis with respect to the coherence times at the different nodes in the same repeater chain. Finally, we present our most extensive result in Section IV.4, where we investigate the optimal placement of quantum repeaters in a 2D plane.
@@ III Model for quantum repeaters
Each pair of neighboring nodes in the chain is connected by an elementary link that can be used to generate elementary-link entanglement. In a chain of $N$ nodes (of which N2N-2 are repeaters), there are N1N-1 elementary links. Generating elementary-link entanglement is a sequence of probabilistic attempts. Each attempt at elementary link $i$ succeeds only with probability $p_{i}$ and takes an amount of time $\Delta t_{i}$. We assume that the attempt duration is limited by the transmission of optical pulses through fiber, giving
@@
[…]
@@
$F_{i}\equiv\braket{\phi^{+}|\rho_{i}|\phi^{+}}=\frac{1+3w_{i}}{4}.$ (8)
@@ IV.2 Optimizing a repeater chain
[…]
Using the repeater-chain model introduced in Section III, we here address the problem of choosing the bright-state parameters for each of the N1N-1 elementary links such that the secret-key rate, defined in Eq. (11), is optimized. A simplified version of this optimization was phrased as a mathematical program by Vardoyan et al. [25] that allows for convex optimization for a specific class of utility functions, a class that was extended further in [97]. Optimization of bright-state parameters under more complicated conditions was performed using genetic algorithms in [34, 38]. Here we follow Vardoyan et al. by making the typical simplifying assumption that each elementary link produces Werner states. In particular, when $\alpha_{i}$ is the bright-state parameter of elementary link $i$, states on that link are created with Werner parameter
@@ IV.4 Repeater placement
In Figure 4, we show the optimal value of $\text{SKR}_{\text{min}}$ that was found for N=0,1,2,,40N=0,1,2,...,40 for fixed $D=300$ km.
Figure 4: Dependence of the optimized network utility ($\text{SKR}_{\text{min}}$, defined in Eq. (19)) on the number of repeaters (NN). End nodes form a square with edges of length $D=300$ km. Shown are the optimization results, a fit function as described by Eq. (22), the minimal number of repeaters ($N_{\text{min}}$, defined in Eq. (21)), the sufficient number of repeaters ($N_{\text{suff, }\delta}$, defined in Eq. (23), with $\delta=0.05$), and the best secret-key rate ($\text{SKR}_{\text{best}}$, defined in Eq. (24)). The error bars in the data are smaller than the marker size and have been suppressed.
@@ V Conclusion
Two open-source Julia packages were developed to make the investigations reported on in this paper possible, QuantumNetworkRecipes.jl [89] and RepeaterPlacement.jl [122]. All the code used for this work has been bundled in a GitHub repository [126], which has been archived at [127].
@@
The data presented in this work has been made available at [128].
@@ Appendix D Simulated annealing
In order to optimize repeater placement as discussed in Section IV.4, we have used simulated annealing. There is a temperature parameter that starts at a high value and is decreased every epoch. Each of the optimizations reported here consists of 25 epochs, with 15 iterations per epoch. For each iteration in epoch i=1,2,,25i=1,2,...,25, the used temperature is $t=t_{0}*0.8^{i}$. The value of $t_{0}$ at $D=300$ km is either 5, 10, or 15, and for other values of $D$, these values are changed in a way that is directly proportional to $D$. As reported in Section IV.4, the temperature is used as the learning rate of the gradient-descent algorithm; this is the reason why we set $t_{0}$ proportional to $D$, as the step size with which the repeaters are displaced should scale with the distances between the end nodes, as otherwise the algorithm will place repeater nodes far outside of the convex hull of the end nodes for small values of $D$ while they will not be varied significantly enough for large values of $D$. Apart from being used as the learning rate, the temperature affects the optimization in the following ways:
@@ Appendix E More repeater-placement results
Figure 10: Dependence of the guaranteed minimal quality of service, $\text{SKR}_{\text{min}}$ as defined in Eq. (19), on NN for $D=50$ km and $D=500$ km. Shown are the optimization results, a fit function as described by Eq. (22), the minimal number of repeaters ($N_{\text{min}}$, defined in Eq. (21)), the sufficient number of repeaters ($N_{\text{suff, }\delta}$, defined in Eq. (23), with $\delta=0.05$), and the best secret-key rate ($\text{SKR}_{\text{best}}$, defined in Eq. (24)). The error bars in the data are smaller than the marker size and have been suppressed.
In this appendix, we show some additional solutions for the repeater-placement optimization discussed in Section IV.4. In Figure 3, we have shown the best repeater placements we have found for $N=4,5,6,7$ and $D=300$ km. Here, in Figure 9, we give some examples of how solutions look for larger values of $N$. We observe that as the number of repeaters grows, the solutions become more irregular and less symmetric. This is likely the case because the optimization problem is more difficult than for a smaller number of repeaters and hence the optimization has not had the chance to converge as well within the 25 epochs of 15 iterations each. Better results would require more computational resources or more efficient optimizations (as discussed in Appendix C, pathfinding is currently the main bottleneck in the optimizations, especially when the number of repeaters is large). Moreover, we observe that as $N$ increases, an increasing number of repeaters is not part of any of the optimal paths between the end nodes. Therefore, the network utility ($\text{SKR}_{\text{min}}$, as defined in Eq. (19)) would not decrease in case those repeaters were removed. This is in line with our observation that the utility stops growing with the number of repeaters for large values of $N$: at some point, adding more repeaters does not allow one to increase the network utility, as introducing them to the paths would reduce the end-to-end fidelity for only a small gain in entangling rate.
In Figure 4 we have shown the dependence of the optimized network utility function $\text{SKR}_{\text{min}}$ on the number of repeaters $N$ for $D=300$ km. Here, in Figure 10, we show what this dependence looks like for a few different values of $D$. For $D=500$ km, while it is clearly visible that the rate at which $\text{SKR}_{\text{min}}$ increases with $N$ slows down for larger values of $N$, it still appears to be growing at the largest value, $N=40$. For this reason, the estimates of $N_{\text{suff},\delta}$ and $\text{SKR}_{\text{best}}$ are perhaps a bit less reliable than those for smaller values of $D$, as a significant part of the function may still be hidden beyond $N=40$. Obtaining significant data for $D>500$ km would probably require solving the optimization for larger values of $N$, which may become computationally expensive.